OpenStudy (anonymous):

If x + 3 / 3 = y + 2 / 2, then x / 3 = ______?

OpenStudy (andriod09):

are they fractions or division?

OpenStudy (anonymous):

Fractions.

OpenStudy (andriod09):

\[\frac {x+3} {3} = \frac {y+2} {2}\] then \[\frac {x} {3} = \frac {w} {z}\] where w and z are the unknown variables you are trying to find out, correct?

OpenStudy (anonymous):

Correct.

OpenStudy (andriod09):

ok then, tell me what you already know about this problem?

OpenStudy (anonymous):

that x / 3 = w / z?

OpenStudy (andriod09):

besides that, what can you tell me about the first part of the problem?

OpenStudy (anonymous):

z = 2??

OpenStudy (andriod09):

not exactly, first, do the top of the fractions of the first part. on the \[\frac {x+3} {3} = \frac {y+2} {2}\] do \[(x+2)(y+2)\]

OpenStudy (anonymous):

this is too confusing.

OpenStudy (andriod09):

do this and tell me what you get: \[(x+3)(y+2)\]

OpenStudy (anonymous):

what am i to do with it? multiply? add?

OpenStudy (andriod09):

distributive property, take the x and multiply it to the 3, the y, and the 2. then do the 3 and multiply that to the y and the 2

OpenStudy (anonymous):

nevermind all this. thanks for helping though.

OpenStudy (andriod09):

meldalz then for trying? :P

OpenStudy (anonymous):

of course.